Steam at 5 bar and 250 Deg is contained in

a rigid vessel of volume 2.2 m, The vessel

is cooled until the steam is just dry saturated.

Calculate the heated rejected, the change in

entropy, the final pressure and the mass of

the steam.



1
Expert's answer
2021-05-25T10:10:14-0400

"\nu=\frac{pV}{RT}=253~mol,"

"Q=\Delta U=3\nu R\Delta T=946~kJ,"

"S=\frac{Q}{\Delta T}=6300~\frac JK,"

"\frac{p_1}{T_1}=\frac{p_2}{T_2}\implies p_2=p_1\frac{T_2}{T_1}=3.6~bar,"

"m=\nu M=4.56~kg."